# Hints-In-Browser: Benchmarking Language Models for Programming Feedback Generation

### v1

@@ 5.3 Results for Quality Performance
For each model and domain, we report aggregated performance over $(T,P_{r})$ instances in the evaluation set. For the program repair scenario, we compute performance from three independent runs and report averaged results as mean (stderr). [^3] For models with fine-tuning, a run means a new model training using fine-tuning pipeline and then inference. For the hints generation scenario, we report performance for only the first run, given the scale of annotations; here, we also report aggregate results across domains. Figure 6 show the results for quality performance and we summarize the main takeaways below.
Effect of Fine-Tuning. We first look at the effectiveness of the fine-tuning pipeline. The performance of base models (Llama-8B and Phi-3.8B) is substantially lower than GPT-4 and on par with GPT-3.5. After fine-tuning, the models’ performance (Llama-8B-dom and Phi-3.8B-dom) surpasses GPT-3.5 on all domains; in particular, Llama-8B-dom’s performance comes close to GPT-4 on IntroPyNUS.
Effect of Web Conversion. An important question is how web conversion and quantization from 16-bit to 4-bit representation affects the quality of feedback generation, as model sizes reduce by a factor of over $3.5$ times. Figure 6 showcases the performance of $4$ models along with their web versions. When looking at aggregated results for all domains on hints quality, we see that Llama-3 models show a smaller drop in performance after web conversion, in contrast to drop for Phi-3 models. [^4] It would useful to compare with the ONNX Runtime engine for Phi-3 web models as mentioned in Section 7.
Effect of Number of Repairs. An important hyperparameter is $n_{r}$, i.e., the total number of program repairs generated by the model as part of the hint generation process. Choosing a higher value can increase hint quality but also increase the inference time. Figure 7 shows results for feedback quality performance with a lower value of $n_{r}=3$, in contrast to $n_{r}=10$ in Figure 6. These results show that $n_{r}=3$ could provide a reasonable quality-time trade-off, especially, on IntroPyNUS.
@@ 6 Evaluation: Web App And Inference Times
Web App. The web app demonstrating the hints-in-browser workflow is accessible at https://hints-in-browser.netlify.app/; see a screenshot in Figure 8. This web app provides a Python programming environment for learners: they can select a programming task in the domain and write/execute programs. This programming environment uses Pyodide [17], a popular browser-based Python interpreter that allows for in-browser Python execution without the need to interact with a server. The feedback generation technique is exposed to learners as a hint button. The model is downloaded and cached locally when the learner accesses the web app for the first time. When a learner requests a hint in the web app, our technique runs on the browser using the WebLLM engine and provides hints to the learner directly in the app. After an initial download and caching of the model, a learner can continue to work on programming tasks and benefit from hints-in-browser without internet connectivity. We will provide the full implementation in the supplementary material.
@@
Results Across Hardware Configurations. Next, we report results regarding total clock time (seconds) for inference during hint generation, model reloading time from the browser cache whenever the app is refreshed, and model downloading time initially. In-browser inference engines require a GPU-equipped machine; hence, we only consider compatible hardware configurations accessible to us. Figure 9 reports results for different hardware configurations, including current price estimates thereof. More concretely, we consider IntroPyNUS domain, set $n_{r}=3$, and report times averaged over the five programming tasks. As a point of comparison, when implementing this technique using OpenAI’s GPT-4 model based on workflow in Figure 1(a), the end-to-end times for feedback generation and serving could be over $30$ seconds [9]. This highlights that the inference times for in-browser inferences are competitive with existing feedback generation workflows for compatible machines (e.g., a GPU-equipped consumer laptop starting at $1200$ USD); machines with more powerful GPUs report a drop in inference times to $20$ seconds for the Phi-3-3.8B model.
Hardware Configuration Llama-3-8B-IntroPyNUS-web Phi-3-3.8B-IntroPyNUS-web
Machine Year Price (USD) Inference (s) Reload (s) Inference (s) Reload (s)
Apple M3 Max 2024 $3500$ $28$ $7$ $20$ $4$
Apple M2 Pro 2023 $2200$ $56$ $9$ $34$ $4$
Apple M1 Pro 2021 $1600$ $71$ $12$ $45$ $6$
Apple M2 Air 2022 $1100$ $129$ $12$ $54$ $10$
Desktop RTX 3080 Ti 2021 $2000$ $38$ $22$ $36$ $11$
Laptop RTX 3060 2022 $1200$ $48$ $25$ $44$ $15$
Laptop RTX 2060 2020 $850$ $624$ $30$ $310$ $14$
Figure 9: Inference times in the web app for $n_{r}=3$. Typical initial download times on a $100$ Mbps connection are around $6$ mins for Llama-3-8B ($4.2$GB file) and $3$ mins for Phi-3-3.8B ($2$GB file).
@@ 7 Concluding Discussions
We investigated language models for programming feedback generation and benchmarked their performance across several criteria crucial for real-world educational deployments. We showed how to leverage the new paradigm of in-browser inference that allows these models to run directly inside the browser and introduced hints-in-browser as a new deployment workflow. As a concrete instance of hints-in-browser, we showcased the efficacy of a fine-tuned Llama-3-8B and Phi-3-3.8B 4-bit quantized model using the WebLLM engine across different performance criteria.
Limitations and Future Work. Next, we discuss some limitations of our current work and ideas for tackling them in the future. First, we experimented with Llama-3-8B and Phi-3-3.8B as base models; it would be interesting to consider variants of Llama-3-8B and Phi-3-3.8B fine-tuned on programming datasets and even smaller models possibly obtained by distilling these base models. Second, we leveraged WebLLM’s in-browser inference engine; it would be useful to compare it with alternatives such as the ONNX Runtime engine. Third, we studied the scenario of programming feedback generation; it would be worth exploring in-browser language models for other domains.

### v2

@@ 5.3 Results for Quality Performance
For each model and domain, we report aggregated performance over $(T,P_{r})$ instances in the evaluation set. For the program repair scenario, we compute performance from three independent runs and report averaged results as mean (stderr). For models with fine-tuning, a run refers to training a new model using our fine-tuning pipeline followed by inference. For the hints generation scenario, we report performance for only the first run, given the scale of annotations; here, we also report aggregate results across domains. Table 1 shows the results for quality performance and we summarize the main takeaways below.
Effect of Fine-Tuning. We first look at the effectiveness of the fine-tuning pipeline. The performance of base models (Llama-8B and Phi-3.8B) is substantially lower than GPT-4 and on par with GPT-3.5. After fine-tuning, the models’ performance (Llama-8B-dom and Phi-3.8B-dom) surpasses GPT-3.5 on all domains; in particular, Llama-8B-dom’s performance surpasses GPT-4o-mini on all domains and comes close to GPT-4 on IntroPyNUS.
Effect of Web Conversion. An important question is how web conversion and quantization from 16-bit to 4-bit representation affects the quality of feedback generation, as model sizes reduce by a factor of over $3.5$ times. Table 1 showcases the performance of $4$ models along with their web versions. When looking at aggregated results for all domains on hints quality, we see that Llama-3 models show a smaller drop in performance after web conversion, in contrast to drop for Phi-3 models.
Effect of Number of Repairs. An important hyperparameter is $n_{r}$, i.e., the total number of program repairs generated by the model as part of the hint generation process. Choosing a higher value can increase hint quality but also increase the inference time. Table 2 shows results for feedback quality performance with a lower value of $n_{r}=3$, in contrast to $n_{r}=10$ in Table 1. These results show that $n_{r}=3$ could provide a reasonable quality-time trade-off, especially, on IntroPyNUS.
Technique Domain:KarelAlgo Domain:KarelAlgo:New
$n_{r}=10$ HGoodRPassREdit HGoodRPassREdit
Llama-3-8B $24$ $41$ $(3)$ $19$ $(1)$ $22$ $37$ $(10)$ $18$ ( $(\ \ 2)$
Llama-3-8B-dom $71$ $90$ $(3)$ $22$ $(0)$ $67$ $74$ ( $(\ \ 7)$ $33$ ( $(\ \ 4)$
Phi-3-3.8B $14$ $33$ $(3)$ $39$ $(6)$ 00 $7$ ( $(\ \ 4)$ $17$ $(10)$
Phi-3-3.8B-dom $38$ $67$ $(3)$ $14$ $(1)$ $33$ $48$ ( $(\ \ 7)$ $21$ ( $(\ \ 2)$
(e) Results on new tasks in a domain.
Dataset (%) Domain:IntroPyNUS
$n_{r}=10$ minEdit RPassHGood%
00 $82$ $(10)$ $93$ $(1)$ $52$
$10$ $24$ ( $(\ \ 4)$ $91$ $(1)$ $60$
$50$ $10$ ( $(\ \ 2)$ $92$ $(0)$ $68$
$100$ $6$ ( $(\ \ 1)$ $100$ $(0)$ $96$
(f) Results with varying dataset sizes.
Figure 6: (a) Results for KarelAlgo and KarelAlgo:New; (b) Results for Llama-3-8B-dom for IntroPyNUS when varying the size of the training dataset.
Generalization for New Tasks in a Domain. In Figure 6, we report additional results on how the fine-tuned models perform on new tasks in the same domain that were unseen during training. Here, we consider the KarelAlgo domain and introduce $3$ new programming tasks in the domain, namely ArithmeticSeriesUptoN, ReverseList and ConcatenateLists. We call this new evaluation set as KarelAlgo:New that also includes $3$ buggy programs for each of the new tasks. We see that the performance boost by fine-tuning is transferred to new problems in the same domain.
Varying the Size of the Training Dataset. In Figure 6, we show additional results where we vary the percentage of the training dataset used for the IntroPyNUS domain. The column labeled minEdit measures the token edit distance between a buggy program in the evaluation set and its closest counterpart in the corresponding training set, reported as mean (stderr).
@@ 6 Evaluation: Web App And Inference Times
Web App. The web app demonstrating the hints-in-browser workflow is accessible at https://hints-in-browser.netlify.app/; see a screenshot in Figure 7. This web app provides a Python programming environment for learners: they can select a programming task in the domain and write/execute programs. This programming environment uses Pyodide [17], a popular browser-based Python interpreter that allows for in-browser Python execution without the need to interact with a server. The feedback generation technique is exposed to learners as a hint button. The model is downloaded and cached locally when the learner accesses the web app for the first time. When a learner requests a hint in the web app, our technique runs on the browser using the WebLLM engine and provides hints to the learner directly in the app. After an initial download and caching of the model, a learner can continue to work on programming tasks and benefit from hints-in-browser without internet connectivity. We provide the full implementation in our Github repository.
@@
Results Across Hardware Configurations. Next, we report results regarding the inference time during hint generation, model reloading time from the browser cache whenever the app is refreshed, and model downloading time initially. [^3] For the web models, typical initial download times on a $100$ Mbps connection are around $6$ mins for Llama-3-8B ($4.2$GB file) and $3$ mins for Phi-3-3.8B ($2$GB file). In-browser inference engines require a GPU-equipped machine; hence, we only consider compatible hardware configurations accessible to us. Table 3 provides a comparison of different deployment workflows for feedback generation, as shown in Figure 1, while considering different hardware configurations. More concretely, we consider the IntroPyNUS domain, set $n_{r}=3$, and report results averaged over five programming tasks. Our results highlight that the inference times for in-browser inferences are competitive with existing feedback generation workflows for compatible machines (e.g., a GPU-equipped consumer laptop starting at $1200$ USD); machines with more powerful GPUs report a drop in inference times to $20$ seconds for the Phi-3-3.8B model. It is worth noting that we could achieve lower inference times by reducing the number of intermediate repairs and potentially by fine-tuning a base model to reduce the number of output tokens generated, e.g., fewer tokens generated as part of the chain-of-thought process. Table 3 further reports on the performance criteria of cost per hint request and privacy aspects of the learner’s data.
Results for Local Models Running Natively. We further evaluated the latencies of a 4-bit quantized Llama-3-8B model running natively on an Apple M1 Pro Laptop, using the same evaluation setup as in Table 3 in two formats: GGUF [37] and MLX [38]. The GGUF format model executed using Ollama had an inference time of about 70 seconds, while the MLX format model (optimized for Apple silicon via the MLX framework) had an inference time of 35 seconds. In comparison, the web version of the model using the WebLLM in-browser inference engine had an inference time of about 70 seconds, thereby closely matching the inference time of the GGUF format model. Notably, in comparison to local models running natively, in-browser inference engines offer significant practical advantages, such as cross-platform compatibility through WebGPU, ease of access without additional setup, and enhanced security via a sandboxed environment.
Table 3: Comparison of different deployment workflows for feedback generation, as shown in Figure 1. Results are reported for the IntroPyNUS domain and $n_{r}=3$. Under the “Hardware Configuration”, L: indicates laptops and D: indicates desktops. Under the “Performance”, the “Inference (s)” and “Cost (USD)” columns report results per hint request, and the “Privacy” column highlights privacy aspects of the learner’s data across different workflows. (We estimated the energy cost of locally invoking the domain-specific web models on a Apple M1 Pro, i.e., L:M1 Pro and it was approximately $6.8\!\!\times\!\!10^{-5}$ USD.)
Model Hardware Configuration Performance
Machine Year Price (USD) Inference (s) Reload (s) Cost (USD) Privacy
GPT-4 n/a n/a n/a $34$ 00 $5.7\!\!\times\!\!10^{-2}$ Fig 1a: External org
GPT-4o mini n/a n/a n/a $15$ 00 $1.0\!\!\times\!\!10^{-3}$ Fig 1a: External org
Llama-3-8B-dom n/a n/a n/a $31$ 00 $2.1\!\!\times\!\!10^{-3}$ Fig 1b: External server
Llama-3-8B-dom-web L:M3 Max 2024 $3500$ $28$ $7$ $0^{\star}$ Fig 1c: User local
L:M2 Pro 2023 $2200$ $56$ $9$
L:M1 Pro 2021 $1600$ $71$ $12$
L:M2 Air 2022 $1100$ $129$ $12$
D:RTX 3080 Ti 2021 $2000$ $38$ $22$
L:RTX 3060 2021 $1200$ $48$ $25$
L:RTX 2060 2021 $850$ $624$ $30$
Phi-3-3.8B-dom-web L:M3 Max 2024 $3500$ $20$ $4$ $0^{\star}$ Fig 1c: User local
L:M2 Pro 2023 $2200$ $34$ $4$
L:M1 Pro 2021 $1600$ $45$ $6$
L:M2 Air 2022 $1100$ $54$ $10$
D:RTX 3080 Ti 2021 $2000$ $36$ $11$
L:RTX 3060 2021 $1200$ $44$ $15$
L:RTX 2060 2021 $850$ $310$ $14$
@@ 7 Concluding Discussions
We investigated language models for programming feedback generation and benchmarked their performance across several criteria crucial for real-world educational deployments. We showed how to leverage the new paradigm of in-browser inference that allows these models to run directly inside the browser and introduced hints-in-browser as a new deployment workflow. As a concrete instance of hints-in-browser, we showcased the efficacy of a fine-tuned Llama-3-8B and Phi-3-3.8B 4-bit quantized models using the WebLLM engine across different performance criteria.
Limitations and Future Work. Next, we discuss some limitations of our current work and ideas for tackling them in the future. First, we experimented with Llama-3-8B and Phi-3-3.8B as base models; it would be interesting to consider variants of Llama-3-8B and Phi-3-3.8B fine-tuned on programming datasets and even smaller models possibly obtained by distilling these base models. In particular, it would be useful to experiment with new small models from the Llama-3.2 family, which also have the potential to perform well on mobile and edge hardware [39]. Second, we leveraged WebLLM’s in-browser inference engine; it would be useful to compare it with alternatives such as the ONNX Runtime engine. Third, we studied the scenario of programming feedback generation; it would be worth exploring in-browser language models for other domains. Last, it would be interesting to examine how this workflow performs in terms of sustainability. While we compared different workflows primarily based on quality and cost, it would also be useful to conduct a systematic evaluation based on energy usage.